# Entanglement entropy of two disjoint intervals and spin structures in interacting chains in and out of equilibrium

### 15 Dec 2023 v1

@@ 4.2 The free boson compactified on a circle
The analytical continuation of (4.5) and, accordingly, the limit $\alpha=1$ are not known in general. However, it is possible to perform the analytic continuation and the limit at some special points and for the small $x$ expansion. The small $x$ expansion for $\eta\neq 1$ reads Calabrese2011Entanglement
@@ 4.3 Thirring model on a Riemann surface
To derive the tripartite information in the XXZ chain in the fermionic picture we need to evaluate the partition functions of the massless Thirring model on a higher genus closed Riemann surface for different spin structures. Several papers Freedman1988Thirring ; Freedman1989Thirring ; Wu1989Determinants ; Sachs1996 have already tackled with this problem, by manipulating fermionic path integrals, and the results display inconsistencies about factors and the sign of the partition functions, as e.g. commented in Ref. Sachs1996 . In this work we fix these ambiguities through entanglement considerations in spin chains. We require that the result for the tripartite information obtained in the fermionic picture reproduces the one obtained in the bosonic picture Calabrese2009Entanglement1 and that for $\lambda=0$ the results reduce to free Dirac fermions. In this way we fix the sign discrepancy, but we do not have access to an overall factor common to all spin structures, which is irrelevant for spin chain considerations.
In this way we conclude that for purely imaginary period matrix $\Omega$ we have
$Z(\vec{\varepsilon},\vec{\delta};\lambda)=(-1)^{4\vec{\varepsilon\cdot\delta}}Z(\vec{0},\vec{0};0)\ f_{\alpha}(\lambda)\ \frac{\Theta\left[\begin{matrix}\vec{E}\\ \vec{F}\end{matrix}\right]\left(\vec{0}|\tilde{\Omega}\right)}{\left[\Theta(\vec{0}|\Omega)\right]^{2}}\;,$ (4.11)
where $Z(\vec{0},\vec{0};0)$ is the partition function for free Dirac fermions with all antiperiodic boundary coniditions, $f_{\alpha}$ are unknown functions of $\lambda$ (neglected in Fig. 1) such that $f_{\alpha}(0)=0$, $\tilde{\Omega}$ is a $2(\alpha-1)\times 2(\alpha-1)$ matrix defined by
@@
carry the information about the spin structure. The formula (4.11) differs from the result of Refs Freedman1988Thirring ; Freedman1989Thirring ; Wu1989Determinants by the sign $(-1)^{4\varepsilon\cdot\delta}$. In the free case $\lambda=0$, studied in Ref. Alvarez-Gaume1986 , the partition function of odd spin structures vanishes so the sign is irrelevant, while the relation between different spin structures,
$\frac{Z(\vec{\varepsilon},\vec{\delta};0)}{Z(\vec{0},\vec{0};0)}=\Bigg{|}\Theta\left[\begin{matrix}\vec{\varepsilon}\\ \vec{\delta}\end{matrix}\right]\left(\vec{0}|\Omega\right)/\Theta\left(\vec{0}|\Omega\right)\Bigg{|}^{2}\,,$ (4.14)
is reproduced by the formula (4.11). In the next section we comment how the formula gives the correct tripartite information for the XXZ spin chain.
@@ 4.4 Spin structures
for some non-universal constant $\mathcal{C_{\alpha}}$, that is independent of the spin structure. The numerator in Eq. (4.15) is expressed in terms of a $2(\alpha-1)\times 2(\alpha-1)$ matrix $\tilde{\Omega}$. It is possible to express it in terms of $\alpha-1\times\alpha-1$ matrices, by resorting to the identity
@@
The tripartite information. (3.18) is obtained by summing over all spin structures. The constant $\mathcal{C}_{\alpha}$ can be fixed by requiring that the tripartite information vanishes at $x\to 0$, which is a consequence of clustering (equivalent trick was used in Ref. Calabrese2011Entanglement for the Ising CFT). This fixes the constant to $\mathcal{C}_{\alpha}=1$. In this way we obtain the result in Eq. (2.10). In Appendix B we show that formula (B.1) reproduces the known result (4.5) obtained from the consideration of the free compactified boson.
@@
The small $x$ expansion of (2.10) can be derived by using the methods developed in Ref. Calabrese2011Quantum for the free compactified boson and Ising CFT. Details of the procedure are presented in Appendix F. We are going to look separately at three different classes of spin structures.
@@
The tripartite information of the XXZ chain is obtained by summing over all spin structures (see Eq. (3.18)). The small $x$-expansion for the tripartite information in the XXZ chain, given by Eq. (4.6), comes from the first term in (4.17) and the second term in (4.19) (and taking the logarithm). Note that the second term in (4.19) is subleading for $\eta>1/2$, but summing over spin structures kills the leading term.
@@ 5.1 Fermionic XXZ chain
In the previous section we have computed the contribution of each spin structure to the tripartite information of the XXZ chain. Having, in particular, the result for the spin structure $\vec{\varepsilon}=\vec{\delta}=\vec{0}$ (antiperiodic boundary conditions on all cycles), we can compute the tripartite information in the fermionic XXZ chain (2.2), where only this spin structure contributes. In this way we obtain the result presented in Eq. (2.4). In Fig. 7 the result is compared to the one for the XXZ spin chain, for $\alpha=2$. The formula (2.4) is manifestly invariant under $\eta\leftrightarrow 1/(4\eta)$, which is a symmetry around the free point $\eta=1/2$. This symmetry should be contrasted to the symmetry $\eta\leftrightarrow 1/\eta$ of the free compactified boson result in Eq. (4.5). At the free point $\eta=1/2$ we get $I_{3,\mathrm{f}}^{(\alpha)}(x)=0$, because of the identity Headrick2013Bose
@@
Vanishing of the tripartite information in the non-interacting case, i.e. in the Kitaev chain at its phase transition, is in agreement with the known result for the free Dirac fermion Casini2005 ; Casini2009Entanglement ; Casini2009reduced_density .
@@ 5.2 Jordan-Wigner deformations
which are also known as “cluster interactions” Raussendorf2001 ; Doherty2009 ; Son2011 ; Giampaolo2015 . Because of the spin structure, such a deformation has now a nontrivial effect. This leads us to Hamiltonian (2.3), that we call “Jordan-Wigner-$\nu$ deformations of the XXZ chain”. In fact, we will see in Sec. 6 that JW deformations characterise the spin structure almost completely.
@@
The model is a sum of fermionic XXZ chain Hamiltonians (see Eq. (2.2)), on sublattices $\nu\mathbb{Z}+j=\{,\ldots,-\nu+j,j,\nu+j,2\nu+j\ldots\}$, for $j=0,1,\ldots,\nu-1$. The Hamiltonians are mutually commuting so they can be diagonalized simultaneously. In particular, all the fermionic correlation functions in the ground state factorize to those involving only fermions on the same sublattice. This also implies the factorization of the function in Eq. (4.15),
@@
where we use the notation $A|_{j}=A\cap(\nu\mathbb{Z}+j)$ for the interesection of a set $A$ and particular sublattice. The factorization (5.6) is exact on the lattice. In the limit of large subsystems we have a dependence on the configuration only through the cross ratio, so
@@
Thus, the contribution of each spin structure to the tripartite information in the Jordan-Wigner deformation of the XXZ chain can be computed from the same quantity in the XXZ chain. By summing over spin structures, we get the tripartite information in the Jordan-Wigner deformation of the XXZ chain,
$I_{3}^{(\alpha)}(x)=\frac{1}{\alpha-1}\log\left(\frac{1}{2^{\alpha-1}}\sum_{\vec{\varepsilon},\vec{\delta}\in\{0,\frac{1}{2}\}^{\alpha-1}}(-1)^{4\vec{\varepsilon}\cdot\vec{\delta}}\left(\mathcal{F}_{f}^{(\mathrm{XXZ})}\left[\begin{smallmatrix}\vec{\varepsilon}\\ \vec{\delta}\end{smallmatrix}\right](x)\right)^{\nu}\right),$ (5.8)
@@
For $\nu=2$ we have shown that the tripartite information is equal to twice the one of the fermionic XXZ chain (see Eq. (2.14)). Therefore the small $x$ expansion can be read from the small $x$ expansion for the latter, given by Eq. (5.2) and Eqs (5.3) for Rényi index $\alpha=2,3,\ldots$ and by Eqs (2.7) and (5.4) for the von Neumann case.
@@
for the von Neumann tripartite information, where the function $s_{\mathrm{vN}}$ is defined in Eq. (2.8). We notice that any time the tripartite information is negative, the leading term is linear in $x$, as in the non-interacting case $\eta=1/2$, also discussed in Ref. Fagotti2012New .
@@ 6 Spin structure from the fusion of models
The tripartite information of the JW deformation has the form
$e^{(\alpha-1)[I_{3}^{(\alpha;\nu)}(x)+\log 2]}=\sum_{n=1}^{M_{\alpha}}g_{\alpha,n}\gamma_{\alpha,n}^{\nu}-\sum_{n=1}^{\tilde{M}_{\alpha}}\tilde{g}_{\alpha,n}(-\gamma_{\alpha,n})^{\nu}\,,$ (6.1)
where each $\gamma$ is a ratio of partition functions and $g_{\alpha,n}$ are positive integers that represent the degeneracy of the terms. If we consider a large enough number of deformations, we can generally express $\{\gamma_{\alpha,n}\}$ in terms of $\{e^{(\alpha-1)[I_{3}^{(\alpha;\nu)}(x)+\log 2]}\}$ by solving polynomial equations. We have performed this procedure for $\alpha=2$ and $\alpha=3$ in Appendix E. The result for $\alpha=2$ is given in Eq. (2.15).
@@
While a priori effective, the method of with JW spin deformations has a rather serious downside: it requires the analysis of models with next-to-next-to-next-to-nearest-neighbor interactions. This can be numerically challenging both because higher range is generally associated with higher numerical complexity and because the typical lengths of a JW $\nu$-deformation is $\nu$ times the typical length of the original model. We can ease these problems by generalising the deformations to the fusion of different models, as also considered in Ref. Fagotti2012New . We can then set an upper bound to the range, for example next-to-nearest-neighbor interactions, and consider a large enough number of combinations of models that allow us to invert the problem.
@@
where the superscripts indicate the model in which $\mathcal{F}_{f}$ is computed. These are the interleaved model (6.2), the XXZ spin chain (2.1) and the XY spin chain (3.34).
The relation (6.3) can be exploited to compute $\mathcal{F}_{f}$ in the original model, i.e. the XXZ spin chain, for any configuration $A,B,C$. For simplicity we focus on the Rényi index $\alpha=2$, corresponding to the torus. To compute $\mathcal{F}_{f}^{\mathrm{XXZ}}\left[\begin{smallmatrix}\varepsilon\\ \delta\end{smallmatrix}\right](A,B,C)$ we can consider the tripartite information in the XXZ chain and in the interleaved model corresponding to three different values of $\gamma,h$ close to or at the circle $\gamma^{2}+h^{2}=1$. One could be $\gamma=1,h=0$, for which $\mathcal{F}_{f}^{\mathrm{XY}}\left[\begin{smallmatrix}\varepsilon\\ \delta\end{smallmatrix}\right](A,B,C)=(-1)^{4\varepsilon\delta}$ irrespectively of the configuration and the others should be taken not too far from the multicritical point $\gamma=0,h=1$ (and not too close either). This will give us a system of linear equations whose solution gives the quantity of interest, i.e. $\mathcal{F}_{f}^{\mathrm{XXZ}}\left[\begin{smallmatrix}\varepsilon\\ \delta\end{smallmatrix}\right](A,B,C)$ for all torus spin structures $\varepsilon,\delta$. We want to use the parameters $h,\gamma$ close to the circle $h^{2}+\gamma^{2}=1$ because only there the dependence of $\mathcal{F}_{f}$ on the configuration is sufficiently varying to give us distinct enough coefficients of the linear problem (see Sec. 3.3).
Let us use the shorthand notation $\mathcal{C}=(A,B,C)$ for the configuration of interest. Let us denote by $\tilde{A},\tilde{B},\tilde{C}$ three adjacent subsystems that are twice the size of $A,B,C$, i.e. $\tilde{A}=\{1,2,\ldots,2|A|\}$, $\tilde{B}=\{2|A|+1,2|A|+2,\ldots,2|A|+2|B|\}$, $\tilde{C}=\{2(|A|+|B|)+1,2(|A|+|B|)+2,\ldots,2(|A|+|B|+|C|)\}$, and let us denote shortly the whole configuration by $\tilde{\mathcal{C}}=(\tilde{A},\tilde{B},\tilde{C})$. Then the interleaved model satisfies $\mathcal{F}_{f}^{\mathrm{XXZ}\oplus\mathrm{XY}}\left[\begin{smallmatrix}\varepsilon\\ \delta\end{smallmatrix}\right](\tilde{\mathcal{C}})=\mathcal{F}_{f}^{\mathrm{XY}}\left[\begin{smallmatrix}\varepsilon\\ \delta\end{smallmatrix}\right](\mathcal{C})\mathcal{F}_{f}^{\mathrm{XXZ}}\left[\begin{smallmatrix}\varepsilon\\ \delta\end{smallmatrix}\right](\mathcal{C})$. Considering the configuration $\mathcal{C}$ in the XXZ spin chain and the configuration $\tilde{\mathcal{C}}$ in interleaved model for $(\gamma_{1},h_{1})=(1,0)$ and another two values $(\gamma_{j},h_{j})$, $j=1,2$, close to the circle $\gamma^{2}+h^{2}=1$, we obtain the system of equations
@@
where by $\mathrm{XY}_{j}$ we denote the quantum XY chain (3.34) with $(\gamma,h)=(\gamma_{j},h_{j})$ and we have also omitted writing the Rényi index $\alpha$ in the tripartite information $I_{3}^{(\alpha)}$. Solving the linear system of equations provides us the quantity of interest.
In practice, we compute the tripartite information with tensor networks. In the interleaved model we take the values of $\gamma_{j},h_{j}$ for $j=2,3$ exactly on the circle $\gamma^{2}+h^{2}=1$. Instead of now computing $\mathcal{F}_{f}^{\mathrm{XY}}$ directly, using analytical formulas (3.37), we do a “renormalization” of the solution. We assume that the ground state, due to its approximate nature arising from the finiteness of the bond dimension can be better approximated with some effective values $\gamma,h$ different from the initial ones, that depend on the bond dimension. We determine these effective $\gamma,h$ by studying the expectation values of some local observables, that can be computed easily with tensor networks. E.g. we use the formulas
@@
that hold on the circle $\gamma^{2}+h^{2}=1$, plug the numerically obtained expectation values to the left hand side and numerically solve the system of integral equations to obtain effective $(\gamma,h)$. Then we use these effective $\gamma,h$ in computing $\mathcal{F}_{f}^{\mathrm{XY}}$. It turns out that this renormalization step is important, as the developed procedure is sensitive to the details of coefficients in the linear system of equations. In Appendix D we demonstrate explicitly the importance of this step, by comparing some results obtained with and without it. The procedure for computing $\mathcal{F}_{f}^{\mathrm{XY}}$ has been developed in Ref. Fagotti2010disjoint , and discussed also in Refs Maric2023Universality ; Maric2023universality2 . For completeness, we present it in Appendix C.

### 11 Mar 2024 v3

@@ 4.2 The free boson compactified on a circle
The analytical continuation of (4.5) and, accordingly, the limit $\alpha\rightarrow 1^{+}$ are not known in general. However, it is possible to perform the limit $\alpha\rightarrow 1^{+}$ for small $x$. The small $x$ expansion for $\eta\neq 1$ reads Calabrese2011Entanglement
@@ 4.3 Thirring model on a Riemann surface
To derive the tripartite information in the XXZ chain in the fermionic picture we need to evaluate the partition functions of the massless Thirring model on a higher genus closed Riemann surface for different spin structures. Several papers Freedman1988Thirring ; Freedman1989Thirring ; Wu1989Determinants ; Sachs1996 have already tackled with this problem, by manipulating fermionic path integrals, and the results display inconsistencies about factors and the sign of the partition functions, as e.g. commented in Ref. Sachs1996 . In this work we fix such ambiguities by requiring that our results reproduce the known predictions for the entanglement entropies of disjoint blocks in the bosonic picture Calabrese2009Entanglement1 . In this way we fix some sign discrepancies but we do not investigate overall factors common to all spin structures, which are irrelevant to our purposes.
We conclude that, for purely imaginary period matrix $\Omega$, we have
$Z(\vec{\varepsilon},\vec{\delta};\lambda)=(-1)^{4\vec{\varepsilon}\cdot\vec{\delta}}Z(\vec{0},\vec{0};0)\ f_{\alpha}(\lambda)\ \frac{\Theta\left[\begin{matrix}\vec{E}\\ \vec{F}\end{matrix}\right]\left(\vec{0}|\tilde{\Omega}\right)}{\left[\Theta(\vec{0}|\Omega)\right]^{2}}\;,$ (4.11)
where $Z(\vec{0},\vec{0};0)$ is the partition function for the free Dirac fermion with all antiperiodic boundary coniditions, $f_{\alpha}$ are unknown functions of $\lambda$ (neglected in Fig. 1) such that $f_{\alpha}(0)=1$, $\tilde{\Omega}$ is a $2(\alpha-1)\times 2(\alpha-1)$ matrix defined by
@@
carry the information about the spin structure. The formula (4.11) differs from the result of Refs Freedman1988Thirring ; Freedman1989Thirring ; Wu1989Determinants in the sign $(-1)^{4\vec{\varepsilon}\cdot\vec{\delta}}$. In the free case $\lambda=0$, studied in Ref. Alvarez-Gaume1986 , the partition function of odd spin structures vanishes so that sign is irrelevant and formula (4.11) captures the relation between different spin structures
$\frac{Z(\vec{\varepsilon},\vec{\delta};0)}{Z(\vec{0},\vec{0};0)}=\Bigg{|}\Theta\left[\begin{matrix}\vec{\varepsilon}\\ \vec{\delta}\end{matrix}\right]\left(\vec{0}|\Omega\right)/\Theta\left(\vec{0}|\Omega\right)\Bigg{|}^{2}\,.$ (4.14)
@@ 4.4 Spin structures
for some non-universal constant $\mathcal{C_{\alpha}}$, that is independent of the spin structure. The numerator in Eq. (4.15) is expressed in terms of a $2(\alpha-1)\times 2(\alpha-1)$ matrix $\tilde{\Omega}$. It is possible to express it in terms of $(\alpha-1)\times(\alpha-1)$ matrices, by resorting to the identity
@@
The tripartite information (3.18) is obtained by summing over all spin structures. The constant $\mathcal{C}_{\alpha}$ can be fixed by requiring that the tripartite information vanishes at $x\to 0$, which is a consequence of clustering (an equivalent trick was used in Ref. Calabrese2011Entanglement for the Ising CFT). This fixes the constant to $\mathcal{C}_{\alpha}=1$. In this way we obtain the result in Eq. (2.10). In Appendix B we show that formula (B.1) reproduces the known result (4.5) obtained from studying the free compactified boson.
@@
The small $x$ expansion of (2.10) can be derived using the methods developed in Ref. Calabrese2011Quantum for the free compactified boson and Ising CFT. Details of the procedure are presented in Appendix F. We are going to look separately at three different classes of spin structures.
@@
The small $x$-expansion for the tripartite information in the XXZ chain, given by Eq. (4.6), comes from the first term in (4.17) and the second term in (4.19) (and taking the logarithm). Note that the second term in (4.19) is subleading for $\eta>1/2$, but summing over spin structures kills the leading term.
@@ 5.1 Fermionic XXZ chain
In the previous section we have computed the contribution of each spin structure to the tripartite information of the XXZ chain. From the result for the spin structure $\vec{\varepsilon}=\vec{\delta}=\vec{0}$ (antiperiodic boundary conditions on all cycles), we can readily compute the tripartite information in the fermionic XXZ chain (2.2). In this way we obtain the result presented in Eq. (2.4). In Fig. 7 the result is compared to the one for the XXZ spin chain with $\alpha=2$. Prediction (2.4) is manifestly invariant under $\eta\leftrightarrow 1/(4\eta)$, which is a symmetry around the free point $\eta=1/2$. This symmetry should be contrasted to the symmetry $\eta\leftrightarrow 1/\eta$ of the free compactified boson in Eq. (4.5). At the free point $\eta=1/2$ we get $I_{3,\mathrm{f}}^{(\alpha)}(x)=0$, because of the identity Headrick2013Bose
@@
The vanishing of the tripartite information in the non-interacting case, i.e. in the Kitaev chain at its phase transition, is in agreement with the known result for the free Dirac fermion Casini2005 ; Casini2009Entanglement ; Casini2009reduced_density .
@@ 5.2 Jordan-Wigner deformations
which are also known as “cluster interactions” Raussendorf2001 ; Doherty2009 ; Son2011 ; Giampaolo2015 . Because of the spin structure, such a deformation has now a nontrivial effect. This leads us to Hamiltonian (2.3), that we call “Jordan-Wigner-$\nu$ deformations of the XXZ chain”. We will see in Sec. 6 that JW deformations characterise the spin structure almost completely.
@@
The model is a sum of fermionic XXZ chain Hamiltonians (see Eq. (2.2)), on sublattices $\nu\mathbb{Z}+j=\{,\ldots,-\nu+j,j,\nu+j,2\nu+j\ldots\}$, for $j=0,1,\ldots,\nu-1$. The Hamiltonians are mutually commuting so they can be diagonalized simultaneously. In particular, all the fermionic correlation functions in the ground state factorize into those involving only fermions on the same sublattice. This also implies the factorization of the function in Eq. (4.15),
@@
where we use the notation $A|_{j}=A\cap(\nu\mathbb{Z}+j)$ for the intersection between set $A$ and sublattice $j$. The factorization (5.6) is exact on the lattice. In the limit of large subsystems we have a dependence on the configuration only through the cross ratio, so
@@
Thus, the contribution of each spin structure to the tripartite information in the Jordan-Wigner deformations is trivially related to the corresponding contribution in the XXZ chain. By summing over spin structures, we get the tripartite information in the Jordan-Wigner deformation of the XXZ chain,
$I_{3}^{(\alpha;\nu)}(x)=\frac{1}{\alpha-1}\log\left(\frac{1}{2^{\alpha-1}}\sum_{\vec{\varepsilon},\vec{\delta}\in\{0,\frac{1}{2}\}^{\alpha-1}}(-1)^{4\vec{\varepsilon}\cdot\vec{\delta}}\left(\mathcal{F}_{f}^{(\mathrm{XXZ})}\left[\begin{smallmatrix}\vec{\varepsilon}\\ \vec{\delta}\end{smallmatrix}\right](x)\right)^{\nu}\right),$ (5.8)
@@
For $\nu=2$ we have shown that the tripartite information is equal to twice the one of the fermionic XXZ chain (see Eq. (2.14)). Therefore its small $x$ expansion can be read from the small $x$ expansion of the latter, given by Eqs (5.2) and (5.3) for Rényi index $\alpha=2,3,\ldots$ and by Eqs (2.7) and (5.4) for the von Neumann case.
@@
for the von Neumann tripartite information, where the function $s_{\mathrm{vN}}$ is defined in Eq. (2.8). We note that, any time the tripartite information is negative, the leading term is linear in $x$, as in the non-interacting case $\eta=1/2$, also discussed in Ref. Fagotti2012New .
@@ 6 Spin structure from the fusion of models
In this section we propose two different procedures for computing the contribution $\mathcal{F}_{f}\left[\begin{smallmatrix}\vec{\varepsilon}\\ \vec{\delta}\end{smallmatrix}\right]$ to the Rényi-$\alpha$ tripartite information in the XXZ chain, with the main focus on the torus, i.e. $\alpha=2$. The first procedure is based on computing the tripartite information in JW deformations for different values of $\nu$. The second procedure is based on computing the tripartite information in two-site shift invariant interleaved models. We mention that our numerical checks for the contributions of different torus spin structures are based on the second procedure, which we find advantageous for numerical implementations.
@@
The tripartite information of the JW deformations has the form given by Eq. (5.8). If we consider a large enough number of deformations $\nu$, we can generally express $\mathcal{F}_{f}^{(\mathrm{XXZ})}\left[\begin{smallmatrix}\vec{\varepsilon}\\ \vec{\delta}\end{smallmatrix}\right](x)$ in terms of $I_{3}^{(\alpha;\nu)}(x)$ by solving polynomial equations, assuming only smoothness as a function of $x$. We have performed this procedure for $\alpha=2$ and $\alpha=3$ in Appendix E. The result for the R-R spin structure for $\alpha=2$ is also given in Eq. (2.15). While a priori effective, the method with JW spin deformations has a rather serious downside: it requires the analysis of models with next-to-next-to-next-to-nearest-neighbor interactions. This can be numerically challenging both because higher range is generally associated with higher numerical complexity and because the typical lengths of a JW $\nu$-deformation is $\nu$ times the typical length of the original model.
@@
We can ease the problems inherent to the procedure with JW deformations by considering the fusion of different models. We can then set an upper bound to the range, for example next-to-nearest-neighbor interactions, and consider a large enough number of combinations of models that allow us to invert the problem.
@@
where the superscripts indicate the model in which $\mathcal{F}_{f}$ is computed, namely, the XXZ spin chain (2.1) and the XY spin chain (3.34).
The relation (6.2) can be exploited to compute $\mathcal{F}_{f}$ in the original model, i.e. the XXZ spin chain, for any configuration $A,B,C$. For simplicity we focus on the Rényi index $\alpha=2$, corresponding to the torus. To compute $\mathcal{F}_{f}^{\mathrm{XXZ}}\left[\begin{smallmatrix}\varepsilon\\ \delta\end{smallmatrix}\right](A,B,C)$ we can consider the tripartite information in the XXZ chain and in the interleaved model corresponding to three different values of $\gamma,h$ close to or at the circle $\gamma^{2}+h^{2}=1$. One could be $\gamma=1,h=0$, for which $\mathcal{F}_{f}^{\mathrm{XY}}\left[\begin{smallmatrix}\varepsilon\\ \delta\end{smallmatrix}\right](A,B,C)=(-1)^{4\varepsilon\delta}$ irrespectively of the configuration and the others should be taken not too far from the multicritical point $\gamma=0,h=1$ (and not too close either). This will give us a system of linear equations whose solution gives the quantity of interest, i.e. $\mathcal{F}_{f}^{\mathrm{XXZ}}\left[\begin{smallmatrix}\varepsilon\\ \delta\end{smallmatrix}\right](A,B,C)$ for all torus spin structures $\varepsilon,\delta$. We want to use the parameters $h,\gamma$ close to the circle $h^{2}+\gamma^{2}=1$ because only there the dependence of $\mathcal{F}_{f}$ on the configuration is sufficiently complex to give us distinct enough coefficients of the linear problem (see Sec. 3.3).
Let us use the shorthand notation $\mathcal{C}=(A,B,C)$ for the configuration of interest. Let us denote by $\tilde{A},\tilde{B},\tilde{C}$ three adjacent subsystems that are twice the size of $A,B,C$, i.e. $\tilde{A}=\{1,2,\ldots,2|A|\}$, $\tilde{B}=\{2|A|+1,2|A|+2,\ldots,2|A|+2|B|\}$, $\tilde{C}=\{2(|A|+|B|)+1,2(|A|+|B|)+2,\ldots,2(|A|+|B|+|C|)\}$, and let us denote the whole configuration by $\tilde{\mathcal{C}}=(\tilde{A},\tilde{B},\tilde{C})$. The interleaved model satisfies $\mathcal{F}_{f}^{\mathrm{XXZ}\oplus\mathrm{XY}}\left[\begin{smallmatrix}\varepsilon\\ \delta\end{smallmatrix}\right](\tilde{\mathcal{C}})=\mathcal{F}_{f}^{\mathrm{XY}}\left[\begin{smallmatrix}\varepsilon\\ \delta\end{smallmatrix}\right](\mathcal{C})\mathcal{F}_{f}^{\mathrm{XXZ}}\left[\begin{smallmatrix}\varepsilon\\ \delta\end{smallmatrix}\right](\mathcal{C})$. Considering the configuration $\mathcal{C}$ in the XXZ spin chain and the configuration $\tilde{\mathcal{C}}$ in interleaved model for $(\gamma_{1},h_{1})=(1,0)$ and other two values $(\gamma_{j},h_{j})$, $j=1,2$, close to the circle $\gamma^{2}+h^{2}=1$, we obtain the system of equations
@@
where by $\mathrm{XY}_{j}$ we denote the quantum XY chain (3.34) with $(\gamma,h)=(\gamma_{j},h_{j})$ and we have also omitted writing the Rényi index $\alpha$ in the tripartite information $I_{3}^{(\alpha)}$. The solution to this linear system of equations unfolds the contribution from each the spin structure in XXZ.
In practice, we compute the tripartite information numerically using tensor network algorithms. In the interleaved model we take the values of $\gamma_{j},h_{j}$ for $j=2,3$ exactly on the circle $\gamma^{2}+h^{2}=1$. The accuracy of the numerical simulations is however crucial since a small discrepancy in the parameters results in huge discrepancies in the $\mathcal{F}$s. Thus, instead of computing $\mathcal{F}_{f}^{\mathrm{XY}}$ directly using analytical formulas (3.37), we ease such sensitivity by renormalising the parameters as follows. We assume that the state is close to the ground state of the XY model for some effective $h$ and $\gamma$ and compare the predictions for the expectation values of two local observables with their numerical values. For instance, we use the identities
@@
which are easily obtained considering that the interleaved model is equivalent to an XY chain when restricted to even sites. We extract the left hand sides of (6.4) from the numerical simulations and invert the system of integral equations for $h$ and $\gamma$. Finally, we use such effective $\gamma,h$ for computing $\mathcal{F}_{f}^{\mathrm{XY}}$. Appendix D demonstrates explicitly the importance of this step by comparing some results obtained with and without it. The procedure for computing $\mathcal{F}_{f}^{\mathrm{XY}}$ has been developed in Ref. Fagotti2010disjoint and discussed also in Refs Maric2023Universality ; Maric2023universality2 . For completeness, it is also reported in Appendix C.